Product Data (10 pts) fivAL Mass: 3.3g Percent Yield: (show calculations) Melting Range: 5-1083C Crystal Color: vello
2 dibenzalacetone acetone benzaldehyde MW- 106.13 g/mol bp-1 78°C d 1.04 g/mL MW 234.30 g/mol mp #110-1110 MW-58.08 g/mol bp 56°C d 0.79 g/mL
Its bs Reaction Setup (3 pts) S. OS Initial quantity of NaOH: S. I Initial quantity of benzaldehyde: mL Initial quantity of acetone: mL Theoretical yield of dibenzalacetone based on benzaldehyde: (show calculations) Theoretical yield of dibenzalacetone based on acetone 9 (show calculations) ce tone Limiting Reagent:
Product Data (10 pts) fivAL Mass: 3.3g Percent Yield: (show calculations) Melting Range: 5-1083C Crystal Color: vello
Answer #1

Mass of benzaldehyde = density x volume = 5.1 ml x 1.04 g/ml = 5.304 g

Moles of Benzaldehyde = mass/molar mass = 5.304 g/106.13 g/mol = 0.04998 mol

Molar ratio of benzaldehyde and dibenzalacetone = 2:1

Theoretical moles of dibenzalacetone = 0.04998 mol/2 = 0.02499 mol

Theoretical yield of dibenzalacetone = Moles x molar mass = 0.02499 mol x 234.30 g/mol = 5.855 gms

Mass of acetone = density x volume = 0.79 g/ml x 1.8 ml = 1.422 g

Moles of acetone = 1.422 g/58.08 g/mol = 0.02448 mol

Molar ratio of acetone and dibenzalacetone = 1:1

Theoretical moles of dibenzalacetone = 0.02448 mol

Theoretical mass of dibenzalacetone = moles x molar mass = 0.02448 mol x 234.30 g/mol = 5.736 gms

From the above theoretical yield based on benzaldehyde and acetone we can observe acetone is limiting reagent

Practical mass of dibenzalacetone = 3.83 g

Percent yield = (practical yield/theoretical yield) x 100 = (3.83 g/5.736 g) x 100 = 66.77 %